# Magnetic Field Line Chaos, Cantori, and Turnstiles in Toroidal Plasmas

### v1

@@ VI.2 Magnetic reconnection when $\vec{B}$ is nearly ideal
VI.3 Limitation on reconnection due to electron inertia
The finite electron mass can cause magnetic reconnection even in the absence of any other non-ideal effect. However, (Boozer, 2025d) proved a remarkable result. When electron inertia is the only non-ideal effect in the evolution of a magnetic field $\vec{B}$, there is a related field that evolves ideally. This field is $\vec{\mathcal{B}}\equiv\vec{B}+\vec{\nabla}\times\left((c/\omega_{pe})^{2}\mu_{0}\vec{j}\right)$ with $\omega_{pe}$ the plasma frequency and $\vec{j}$ the current density. In three dimensional space, the practical importance of the ideal evolution of $\vec{\mathcal{B}}$ on its reconnection appears limited. Section VI.2 shows that when the evolution velocity of modified field is chaotic, $\vec{\mathcal{B}}$ will reconnect on a timescale that depends only logarithmically on any non-ideal effect that is diffusive, such as resistivity.
VI.4 Magnetic surface breakup and electrostatic instabilities
(Nevins et al., 2011), (Connor, Hastie & Zocco, 2013), and (Terry et al., 2015) have discussed magnetic surface breakup produced by what called electrostatic instabilities, primarily the ion temperature gradient (ITG) instabilities. The focus has been on the enhancement of the electron heat transport, which is not large when the plasma $\beta\equiv 2\mu_{0}p/B^{2}$ is small.
An effect that can arise before the electron heat transport is appreciable is a modification of the radial electric field required to preserve quasi-neutrality in stellarators. This modification is of great practical importance because neoclassical transport in stellarators usually gives more rapid ion than electron transport, which requires an electric field that confines ions—particularly high charge state impurities. (Helander & Simakov, 2008) discussed the lack of effect of electrostatic microturbulence on the large scale radial electric field though the turbulent Reynolds stress can produce zonal flows that tend reduce the microturbulent transport. They did not discuss the effect of the breaking of the magnetic surfaces by the turbulence. The enhanced radial transport of electrons does modify the radial electric field. The required momentum transport is a subtlety, which can be viewed as an effective perpendicular viscosity force. (Alcusón et al., 2023) found a large expulsion of impurities in W7-X experiments when the microturbulent transport exceeded the expected neoclassical, which requires an explanation.
@@ VI.4 Magnetic surface breakup and electrostatic instabilities
When the phase velocity of $\tilde{v}_{x}$ along $z$ equals the Alfvén speed, $\tilde{u}_{x}=\tilde{v}_{x}$, and the whole motion of the plasma is due to the motion of the magnetic field lines, which means an Alfvén wave. When the phase velocity along $z$ is approximately the ion thermal speed $V_{i}$, then $\tilde{u}_{x}=(V_{i}^{2}/V_{A}^{2})\tilde{v}_{x}$. The ratio $V_{i}^{2}/V_{A}^{2}\sim\beta$, and the magnetic field lines move little compared to the plasma motion. Nevertheless, the velocity of the magnetic field lines is turbulent and, therefore, chaotic, which implies magnetic reconnection that breaks the magnetic surfaces will quickly occur, regardless of how small non-ideal effects represented by the loop voltage may be.
VII Discussion
@@
Once the concepts of magnetic field line chaos, cantori, and turnstiles are understood their importance to determining the behavior of toroidal fusion plasmas becomes clear. This section gives a few examples of their applications, but also expresses concern about the future of innovation within U. S. fusion program.
VII.1 Tokamak disruptions
@@
As noted in the Introduction, the current density required to create an exponentiation $\sigma$ is only linearly dependent of $\sigma$, see Appendix A, as is the timescale for large scale reconnection, see Section VI. Section IV showed an ideal perturbation can lead to chaotic field lines through a large volume of the plasma even though perfect magnetic surfaces are preserved. Section VI shows that time required before reconnection occurs is essentially determined by timescale on which the exponentiation of the chaos grows. The time required for a large scale reconnection has only a logarithmic dependence on resistivity, which makes this time essentially independent of resistivity when the resistivity is small. Shortly after the paper (Boozer, 2022) was was published showing this, (Jardin et al., 2022) carried out simulations of disruptions that were extremely fast due to an ideal mode becoming unstable. The simulations confirmed that the time to disruption was essentially independent of the plasma resistivity to the extent the code could maintain adequate resolution. That is, any large scale mode, resistive or ideal, can produce a disruption on a timescale determined by the growth rate of the mode unless it self-stabilizes at a sufficiently small amplitude. Non-self-stabilizing modes could cause a stellarator to disrupt, but such modes are easier to avoid in stellarators than in tokamaks.
@@
When the magnetic evolution is near ideal, as in a high-temperature tokamak, (Boozer, 2025b) showed only a small fraction of the energy released by the reconnection is directly dissipated—most goes into Alfvén waves. This is unlike reconnection in two-coordinate systems in which the energy released by the reconnection is damped on the same timescale as the reconnection. (Boozer, 2020) showed that shear Alvén waves moving along the reconnecting field lines would cause the flattening of the $j_{||}/B$ profile, which leads to the current spike observed in disruptions. The simulations of (Nardon et al., 2023) were consistent with the predicted timescale, and they saw magnetic fluctuations that appeared to be shear Alfvén waves. When shear Alfvén waves propagate along chaotic field lines they are quickly damped. (Huang & Bhattacharjee, 2022) defined reconnection, even in three-coordinate problem, as the transfer of energy from the reconnecting field to the plasma, not the breaking of field line connections. Their definition, which is common in two-coordinate models, led them to the odd conclusion that the chaos of the field lines was irrelevant to reconnection—only the formation of localized currents, which are caused by Alfvén wave damping, are relevant.
@@
The rapid spreading of impurities across the entire plasma is a commonly observed feature of tokamak disruptions. The impurities greatly reduce the damage produced by disruptions by radiating the plasma thermal energy relatively uniformly over the walls rather than having it concentrated in spots where turnstiles formed within the plasma strike the walls as is often the case with runaway electrons. The reason for the impurity spreading did not have firm theoretical basis, but the Bohm-like transport produced when there is an electron pressure gradient along chaotic field lines, Boozer (2024b) provides and explanation. The simulations of (Nardon et al., 2023) saw electric fields consistent in magnitude, but the equation in their code which gave the electric field was not simply related to the two-fluid equations that Boozer (2024b) used to obtain the Bohm-like transport.
@@
(Breizman et al., 2019) reviewed the problem of the conversion of the plasma current from being carried by near-thermal to being carried by relativistic electrons, which is often viewed as the most dangerous feature of disruptions. Even though the total energy in the relativistic electrons is only about ten percent of the pre-disruption plasma thermal energy, impurity radiation spreads the thermal energy losses almost uniformly over the wall. Relativistic electron losses onto the walls can be extremely concentrated in both space and time, which means they are far more destructive. As pointed out by (Boozer & Pujabi, 2016) the spatial and temporal concentration can be explained by the breaking of the outermost magnetic surface that is confining the runaway electrons in internal bounded chaotic region. As the outermost surface breaks it forms a cantorus. The slower the turnstiles form compared to the timescale for the relativistic electrons to cover the chaotic region the more tightly collimated are the turnstiles on the cantorus. When the turnstiles form rapidly compared to the timescale for the relativistic electrons to cover the chaotic region, the turnstiles become arbitrary large, which spreads the relativistic electrons so broadly on the walls that they do no damage. The highly damaging case is often observed and brings fear, but (Reux et al., 2021) and (Paz-Soldan et al., 2021) have observed on the JET and the DIII-D tokamaks the benign spreading of the relativistic electrons when the plasma becomes unstable to a rapidly growing magnetic perturbation.
@@
(Boozer, 2025c) has pointed out that the frequency with which disruptions occur in tokamaks may be in large part be due to the far larger poloidal flux produced by the plasma current than the change in the poloidal flux that occurs if the profile of the density of the net plasma current $j_{||}$ is varied over the full range of profiles stable to tearing modes—approximately by a factor of six. Equation (2), $\partial\psi_{p}/\partial t=V_{\ell}$, implies the plasma will disrupt within a shorter time than the time required to shutdown the plasma or the natural length of a pulse length unless a spatially constant loop voltage remains consistent with tearing mode stability. The loop voltage is approximately given by $V_{\ell}=2\pi R(\eta j_{||}-j_{bs}-j_{cd})$, where $R$ is the major radius, $\eta$ is the resistivity, which primarily scales as $1/T_{e}^{3/2}$, $j_{bs}$ is the bootstrap current density, and $j_{cd}$ is the density of the externally driven current. The spatial constancy of $V_{\ell}$ is essentially a constraint on the electron temperature profile to avoid disruption produced by tearing modes. The electron temperature profile is determined by microturbulent transport together with the heating and the cooling. In the shutdown of a fusion plasma, the heating source changes radically as the fusion power turns off. The profile of $j_{||}$ can be controlled by plasma heating and current drive profiles. During the flattop of a fusion pulse, the required power to produce a large change in $j_{||}$ is comparable to the power produced by the alpha-particle heating. This is obvious for direct heating, (Boozer, 1988) showed the power required to drive the full plasma current in a power plant is comparable to the power produced by alpha-particle heating. Careful control of the injected power is required to properly control the profile of $j_{||}$, but the available diagnostics may not provide adequate information for such control even if it were in principle possible. It is unclear whether the ratio of the total plasma flux produced by the plasma current versus the change in the poloidal flux if the current profile evolved from disruption stable to unstable has been calculated for the ARC designs of Commonwealth Fusion Systems or what the strategy is for ensuring a stable profile evolution. This disruption issue can be circumvented by the use of stellarator magnetic field to produce the rotational transform by external coil currents rather than by the plasma current.
VII.2 Stellarator physics
@@
Stellarators have a unique external control over the plasma through the coils, which (Boozer, 2004) showed could efficiently produce approximately fifty independent magnetic field distributions, an order of magnitude more than can be produced in axisymmetry. By their nature, stellarators provide steady-state plasma confinement, low recirculating power, as well as robustness against disruption and the loss of positional control within the plasma chamber.
@@
Adequate stellarator confinement requires optimization of the external field. In 1980 stellarators were known to have many attractive properties for fusion power plant, but it was thought that this was precluded by the extremely large neoclassical transport caused by their toroidal asymmetry. (Boozer, 1980) gave a new formulation of the drift equations and a new coordinate system that allowed a relatively fast optimization to reduce the neoclassical transport, and (Nürenberg & Zille, 1988) used this formalism to show that such magnetic configurations exist. The result of modern optimization for a fusion power plant is illustrated by the Infinity Two fusion pilot plant design (Hegna et al., 2025).
@@
(Boozer, 2024a) gave a conceptual design for stellarators with enhanced tritium confinement and edge radiation control. This design used special features of the stellarator to greatly enhance the tritium burn-up fraction, a non-resonant divertor to allow impurity radiation at the edge to spread the thermal power exhaust broadly over the wall, and flattop density and temperature profiles to enhance the fusion power production and reduce the impurity confinement. The low tritium burn-up fraction is is a major issue in obtaining tritium self-sufficiency in toroidal fusion power plants. It should be noted that if a fission source of tritium could be identified that allowed a large fusion to fission power ratio and with a cost of only millions of dollars per kilogram, fusion power plants could be greatly cheapened and simplified by reducing or eliminating their tritium-breeding blankets. A large fusion/fission power ratio would allow the fission tritium producers to be located at secure sites, such as nuclear weapons facilities. Poor confinement in the central part of the stellarator would flatten all profiles, including those of impurities, and as discussed in Section VI.4 the magnetic field chaos produced at non-zero plasma pressure by what is called electrostatic microturbulence could expel impurities.
VII.3 Program priorities
@@
The Merriam-Webster dictionary gives two definitions of innovation: (1) a new idea, method, or device and (2) the introduction of something new. By this definition, innovations cannot be explicitly predicted. Nevertheless, they are a central element in the achievement or practical fusion energy on the shortest possible timescale as well as advancing other important applications of plasma physics, such as space weather. Admiral Rickover led the development of the nuclear submarine, the Nautilus, which began operations approximately fifteen years after the splitting of the uranium nucleus in Berlin. His insights can provide guidance for the rapid achievement of fusion: “All new ideas begin in a non-conforming mind that questions some tenet of the conventional wisdom.” “We shrivel creativity by endless frustrations.” “Good ideas are not adopted automatically. They must be driven into practice with courageous impatience. Once implemented they can be easily overturned or subverted through apathy or lack of follow-up, so a continuous effort is required.” Rickover did not seek to eliminate the persons who were constantly developing new ideas from the program because a precise time line for developing new ideas could not be developed.
The (U.S. Department of Energy, 2025) released the Fusion Science and Technology Roadmap, which prominently featured “Innovate and advance the science and engineering of fusion with well-defined milestones and metrics” on page 7. The inconsistency of innovations “with well-defined milestones and metrics” could be viewed as a wording issue but, whether intended or not, seemingly defines policy. The reviewers of my proposal on magnetic recognized the “various excellent contributions to reconnection research” that I had made but felt that continuing innovations without milestones was inconsistent with funding, a position accepted by the Department of Energy. The recent papers on reconnection (Boozer, 2025c) and (Boozer, 2025d) had no external support. My individual grants have expired and the renewal proposals have not yet been approved. They too may fail if past innovations are taken as having no importance for future funding. The issue is not me, I can retire at any time. But, the success of fusion program depends on innovations. Assigning a low funding priority to persons who innovate relative to those who run established codes to meet “well-defined milestones and metrics” seems inconsistent with a dynamic and successful program.

### v2

@@ 6.2 Magnetic reconnection when $\vec{B}$ is nearly ideal
6.3 Limitation on reconnection due to electron inertia
The finite electron mass can cause magnetic reconnection even in the absence of any other non-ideal effect. However, (Boozer:2025:Electron) proved a remarkable result. When electron inertia is the only non-ideal effect in the evolution of a magnetic field $\vec{B}$, there is a related field that evolves ideally. This field is $\vec{\mathcal{B}}\equiv\vec{B}+\vec{\nabla}\times\left((c/\omega_{pe})^{2}\mu_{0}\vec{j}\right)$ with $\omega_{pe}$ the plasma frequency and $\vec{j}$ the current density. In three dimensional space, the practical importance of the ideal evolution of $\vec{\mathcal{B}}$ on its reconnection appears limited. Section 6.2 shows that when the evolution velocity of modified field is chaotic, $\vec{\mathcal{B}}$ will reconnect on a timescale that depends only logarithmically on any non-ideal effect that is diffusive, such as resistivity.
6.4 Magnetic surface breakup and electrostatic instabilities
(Nevins:2011), (Connor:2013), and (Terry:2015) have discussed magnetic surface breakup that is produced by what are called electrostatic instabilities, primarily the ion temperature gradient (ITG) instabilities. The focus has been on the enhancement of the electron heat transport, which is not large when the plasma $\beta\equiv 2\mu_{0}p/B^{2}$ is small.
An effect that can arise before the electron heat transport is appreciable is a modification of the radial electric field required to preserve quasi-neutrality in stellarators. This modification is of great practical importance because neoclassical transport in stellarators usually gives more rapid ion than electron transport, which requires an electric field that confines ions—particularly high charge state impurities. (Helander:2008) discussed the lack of effect of electrostatic microturbulence on the large scale radial electric field though the turbulent Reynolds stress can produce zonal flows that tend reduce the microturbulent transport. They did not discuss the effect of the breaking of the magnetic surfaces by the turbulence. The enhanced radial transport of electrons due to broken surfaces does modify the radial electric field, but the required momentum transport to break the surfaces is subtle. As the author has discussed but not yet published, the radial transport of the electrons gives an effective perpendicular viscosity force. (Alcison:2023) found a large expulsion of impurities in W7-X experiments when the microturbulent transport exceeded the expected neoclassical, which requires an explanation.
@@ 6.4 Magnetic surface breakup and electrostatic instabilities
When the phase velocity of $\tilde{v}_{x}$ along $z$ equals $V_{A}$, the Alfvén speed, $\tilde{u}_{x}=\tilde{v}_{x}$, and the whole motion of the plasma is due to the motion of the magnetic field lines, which means an Alfvén wave. When the phase velocity along $z$ is approximately the ion thermal speed $V_{i}$, then $\tilde{u}_{x}=(V_{i}^{2}/V_{A}^{2})\tilde{v}_{x}$. The ratio $V_{i}^{2}/V_{A}^{2}\sim\beta$, and the magnetic field lines move little compared to the plasma motion. Nevertheless, the velocity of the magnetic field lines is turbulent and, therefore, chaotic, which implies magnetic reconnection that breaks the magnetic surfaces will quickly occur, regardless of how small non-ideal effects represented by the loop voltage may be.
7 Discussion
@@
Once the concepts of magnetic field line chaos, cantori, and turnstiles are understood their importance to determining the behavior of toroidal fusion plasmas becomes clear. This section gives a few additional examples of their applications, but also expresses concern about the future of innovation within U. S. fusion program.
7.1 Tokamak disruptions
@@
As noted in the Introduction, the current density required to create an exponentiation $\sigma$ is only linearly dependent of $\sigma$, see Appendix A.3, as is the timescale for large scale reconnection, see Section 6. Section 4 showed an ideal perturbation can lead to chaotic field lines through a large volume of the plasma even though perfect magnetic surfaces are preserved. Section 6 shows that time required before reconnection occurs is essentially determined by timescale on which the exponentiation of the chaos grows. The time required for a large scale reconnection has only a logarithmic dependence on resistivity, which makes this time essentially independent of resistivity when the resistivity is small. Shortly after the paper (Boozer:2022) was was published showing this, (Jardin:2022) carried out simulations of disruptions that were extremely fast due to an ideal mode becoming unstable. Their simulations confirmed that the time to disruption was essentially independent of the plasma resistivity to the extent the code could maintain adequate resolution. That is, any large scale mode, resistive or ideal, can produce a disruption on a timescale determined by the growth rate of the mode unless it self-stabilizes at a sufficiently small amplitude. Non-self-stabilizing modes could cause a stellarator to disrupt, but such modes are easier to avoid in stellarators than in tokamaks.
@@
When the magnetic evolution is near ideal, as in a high-temperature tokamak, (Boozer:2025:Magnetic) showed only a small fraction of the energy released by the reconnection is directly dissipated—most goes into Alfvén waves. This is unlike reconnection in two-coordinate systems in which the energy released by the reconnection is damped on the same timescale as the reconnection. (Boozer:2020) showed that shear Alvén waves moving along the reconnecting field lines would cause the flattening of the $j_{||}/B$ profile, which leads to the current spike observed in disruptions. The simulations of (Nardon:2023) were consistent with the predicted timescale, and they saw magnetic fluctuations that appeared to be shear Alfvén waves.
(Heyvaerts:1983) and (Similon:1989) showed that when shear Alfvén waves propagate along chaotic field lines they are quickly damped. (Huang-Bhattacharjee:2022) defined reconnection, even in three-coordinate problem, as the transfer of energy from the reconnecting field to the plasma, not the breaking of field line connections. Their definition, which is common in two-coordinate models, led them to the odd conclusion that the chaos of the field lines was irrelevant to reconnection—only the formation of localized currents, which are caused by Alfvén wave damping, are relevant.
@@
The rapid spreading of impurities across the entire plasma is a commonly observed feature of tokamak disruptions. The impurities greatly reduce the damage produced by disruptions by radiating the plasma thermal energy relatively uniformly over the walls rather than having it concentrated in spots where turnstiles formed within the plasma strike the walls as is often the case with runaway electrons. The reason for the impurity spreading did not have a firm theoretical basis, but the Bohm-like transport that is produced when there is an electron pressure gradient along chaotic field lines, (Boozer:2024:Electric) provides an explanation. The simulations of (Nardon:2023) saw electric fields consistent in magnitude, but the equation in their code which gave the electric field was not simply related to the two-fluid equations that (Boozer:2024:Electric) used to obtain the Bohm-like transport.
@@
(Breizman:2019) reviewed the problem of the conversion of the plasma current from being carried by near-thermal to being carried by relativistic electrons, which is often viewed as the most dangerous feature of disruptions. Even though the total energy in the relativistic electrons is only about ten percent of the pre-disruption plasma thermal energy, impurity radiation spreads the thermal energy losses almost uniformly over the wall. Relativistic electron losses onto the walls can be extremely concentrated in both space and time, which means they are far more destructive. As pointed out by (Boozer-Punjabi:2016) the spatial and temporal concentration can be explained by the breaking of the outermost magnetic surface that is confining the runaway electrons in internal bounded chaotic region. As the outermost surface breaks it forms a cantorus. The slower the turnstiles form compared to the timescale for the relativistic electrons to cover the chaotic region the more tightly collimated are the turnstiles on the cantorus. When the turnstiles form rapidly compared to the timescale for the relativistic electrons to cover the chaotic region, the turnstiles become arbitrary large, which spreads the relativistic electrons so broadly on the walls that they do no damage. The highly damaging case is often observed and brings fear, but (Reux:2021) and (Paz-Soldan:2021) have observed on the JET and the DIII-D tokamaks the benign spreading of the relativistic electrons when the plasma becomes unstable to a rapidly growing magnetic perturbation.
(Boozer:2025:Constraints) has pointed out that the frequency with which disruptions occur in tokamaks may be in large part be due to the far larger poloidal flux produced by the plasma current than the change in the poloidal flux that occurs if the profile of the density of the net plasma current $j_{||}$ is varied over the full range of profiles stable to tearing modes. This flux ratio is approximately a factor of six. Equation (2), $\partial\psi_{p}/\partial t=V_{\ell}$, implies the plasma will disrupt within a shorter time than the time required to shutdown by removing the plasma-produced poloidal flux or the natural length of a flattop pulse unless a spatially constant loop voltage remains consistent with tearing mode stability. As shown in Equation 19, the magnetic field can be held fixed when $\vec{\nabla}V_{\ell}=0$. An implication is that poloidal flux can be removed or maintained without producing a disruption when
$j_{||}=\frac{V_{\ell}}{2\pi R\eta}+j_{bs}+j_{cd}$ (22)
remains in a non-disruptive profile with $\vec{\nabla}V_{\ell}=0$. $R$ is the major radius, $\eta$ is the resistivity, which approximately scales as $1/T_{e}^{3/2}$ but also depends on the charge state of impurities, $Z_{eff}$. The bootstrap current density is $j_{bs}$, and the density of the externally driven current is $j_{cd}$
The spatial constancy of $V_{\ell}$ is approximately a constraint on the electron temperature profile to avoid disruptions produced by tearing modes. The electron temperature profile is determined by microturbulent transport together with the heating and the cooling. In the shutdown of a fusion plasma, the heating source changes radically as the fusion power turns off.
@@
The profile of $j_{||}$ can in principle be controlled by plasma heating and current drive. During the flattop of a fusion pulse, the required power to produce a large change in the $j_{||}$ profile is comparable to the power produced by the alpha-particle heating. This is obvious for direct heating. (Boozer:1988) showed the power required to drive the full plasma current in a power plant is greater than $j_{cd}E_{ch}$ integrated over the plasma volume where $E_{ch}$ is the Connor-Hastie electric field (Connor:1975). The implication is the power required to drive the total plasma current is comparable to the power produced by alpha-particle heating. Careful control of the injected power is required to properly control the profile of $j_{||}$, but the available diagnostics may not provide adequate information for such control even if it were in principle possible. It is unclear whether the ratio of the total plasma flux produced by the plasma current versus the change in the poloidal flux if the current profile evolved from disruption stable to unstable has been calculated for the ARC designs of Commonwealth Fusion Systems or what the strategy is for ensuring a stable profile evolution. This disruption issue can be circumvented by the use of stellarator magnetic field to produce the rotational transform by external coil currents rather than by the plasma current.
7.2 Stellarator physics
@@
Stellarators have a unique external control over the plasma through the coils, which (Boozer:2004) showed could efficiently produce approximately fifty independent magnetic field distributions, an order of magnitude more than can be produced in axisymmetry. By their nature, stellarators provide steady-state plasma confinement, low recirculating power, as well as robustness against disruption and the loss of positional control within the plasma chamber.
@@
Adequate stellarator confinement requires optimization of the external field. In 1980 stellarators were known to have many attractive properties for fusion power plant, but it was thought that this was precluded by the extremely large neoclassical transport caused by their toroidal asymmetry. (Boozer:1980) gave a new formulation of the drift equations and a new coordinate system that allowed a relatively fast optimization to reduce the neoclassical transport. (Nurenberg:1988) used this formalism to show that such magnetic configurations exist. The result of modern optimization for a fusion power plant is illustrated by the Infinity Two fusion pilot plant design (Hegna:2025).
@@
(Boozer:2024:Stellarators) gave a conceptual design for stellarators with enhanced tritium confinement and edge radiation control. This design used special features of the stellarator to greatly enhance the tritium burn-up fraction, a non-resonant divertor to allow impurity radiation at the edge to spread the thermal power exhaust broadly over the wall, and flattop density and temperature profiles to enhance the fusion power production and reduce the impurity confinement. The low tritium burn-up fraction is a major issue in obtaining tritium self-sufficiency in toroidal fusion power plants. It should be noted that if a fission source of tritium could be identified that allowed a large fusion to fission power ratio and with a cost of only millions of dollars per kilogram, fusion power plants could be greatly cheapened and simplified by reducing or eliminating their tritium-breeding blankets. A large fusion/fission power ratio would allow the fission tritium producers to be located at secure sites, such as nuclear weapons facilities. Poor confinement in the central part of the stellarator would flatten all profiles, including those of impurities, and as discussed in Section 6.4 the magnetic field chaos produced at non-zero plasma pressure by what is called electrostatic microturbulence could expel impurities.
7.3 Program priorities
@@
The Merriam-Webster dictionary gives two definitions of innovation: (1) a new idea, method, or device and (2) the introduction of something new. By either definition, innovations cannot be explicitly predicted. Nevertheless, they are a central element in the achievement of practical fusion energy on the shortest possible timescale as well as advancing other important applications of plasma physics, such as space weather. Admiral Rickover led the development of the nuclear submarine, the Nautilus, which began operations approximately fifteen years after the splitting of the uranium nucleus in Berlin. His insights can provide guidance for the rapid achievement of fusion: “All new ideas begin in a non-conforming mind that questions some tenet of the conventional wisdom.” “We shrivel creativity by endless frustrations." “Good ideas are not adopted automatically. They must be driven into practice with courageous impatience. Once implemented they can be easily overturned or subverted through apathy or lack of follow-up, so a continuous effort is required." Rickover did not seek to eliminate the persons who were constantly developing new ideas from the program because a precise time line for developing new ideas could not be given.
The United States Department of Energy (US-DoE:2025) released the Fusion Science and Technology Roadmap, which prominently featured “Innovate and advance the science and engineering of fusion with well-defined milestones and metrics" on page 7. The inconsistency of innovations “with well-defined milestones and metrics" could be viewed as a wording issue but, whether intended or not, seemingly defines policy.
The reviewers of my April 2025 proposal on magnetic reconnection recognized the “various excellent contributions to reconnection research" that I had made but felt that continuing innovations without well-defined milestones, by the instructions they had received, only justified a low funding priority. This position was accepted by the Department of Energy. Recent papers on reconnection (Boozer:2025:Constraints) and (Boozer:2025:Electron) had no external support.
My proposal seeking continued support for “Non-axisymmetric magnetic fields in tokamaks” was submitted on 4 January 2024 and declined on 1 December 2025 for not having clear milestones despite very positive reviews. During the twenty-three months the renewal was under review, I had six publications relevant to the proposal. Four have been discussed in this article: (Boozer:2024:Electric), (Boozer:2025:Constraints), (Boozer:2025:Efficient), and (Boozer:2025:Magnetic). Two others are this article, and (Boozer:2024:Use), which provided a method for studying the unexplored design freedom of tokamak and stellarator coils. Eight additional publications during this period were on stellarators and the article (Boozer:2025:Electron). Could any of these contributions been foreseen for the three to five years needed to obtain grant approval when “well-defined milestones and metrics” are required?
My individual grants have expired, and only one renewal proposal remains. It is on stellarators and will also fail if past innovations are taken as having no importance for future funding. The issue is not just about me. Many group leaders as well as researchers feel too entangled in the spider web of milestones and metrics to do what they know is of greatest importance or to speak out. My situation makes it incumbent for me to try change a very destructive Department of Energy policy. If I am forced to retire, my taxable income will increase by fifty percent. This and my continuing, impactful, and well documented contributions allow me to publicly seek a change. My longterm contributions and their programmatic impact are documented in part at “Boozer-Google Scholar” with many publications single author.
Does anyone actually think that the achievement of fusion energy is made faster and less costly by removing grant support from persons whose research is focused on innovations that satisfy the standard meaning of being too novel to predict?